# SONICS: Synthetic Or Not - Identifying Counterfeit Songs

### v2

@@ Appendix B Implementation Details
Appendix C Human-AI Benchmark
To evaluate the human performance of synthetic song detection, a Huggingface space is titled "Song Arena" (shown in Fig. 9) is created where users can evaluate whether a random song is synthetic or not from a small subset (520 samples in total) of the proposed dataset. A leaderboard (shown in Fig. 10) consisting of the detection rate of the samples belonging to each algorithm as well as generation methods (mostly fake, half fake and full fake songs - as outlined in Section 3.1.2) are also part of the space. Sensitivity (True Positive Rate) is used to evaluate the performance of synthetic song generation algorithms.
Appendix D Prompt Engineering
@@
Instructions: Analyze the provided song lyrics and extract the following elements: 1. Subject Matter: Write what the song is about by providing a summary of the story, narrative, central topic, or events discussed in the song lyrics. 2. Theme: Identify the main theme or message (emotion, life experience, social commentary, or philosophical concept). 3. Target Audience: Define the intended audience (age group, cultural background, or specific interests). 4. Narrative: The story and point of view (e.g., first person or second person). 5. Character Analysis: Main characters (e.g., Protagonist, Antagonist) and their traits. 6. Song Structure: Outline the structure (number of verses, choruses, bridges, intros, outros) and note any unique elements or deviations. 7. Mood: Describe the overall mood (upbeat, melancholic, introspective, etc.). 8. Reference: Identify any cultural, social, time, place, or contextual references. How to respond: You should provide your answer below after the “Answer” section. You are not allowed to use any text formatting (bold, italic, etc.) and narrative (‘Here is the answer’, ‘Below is the response’, etc.) in your answer. Only answer using the following format: subject_matter: “…..” theme: “….” target_audience: “….” narrative: “….” character_analysis: “….” song_structure: “….” mood: “….” reference: “….” Lyrics: {lyrics} Answer:
Table 7: Prompt template for extracting lyrics features (e.g subject matter, theme, mood) from real songs.
@@
You should write the song with metatags following the song structure from "Lyrics Features". In some very rare cases, you can also scarcely include ad libs, or non-lexical vocables.
@@
Table 8: Prompt template for generating song lyrics from lyrics features.
@@
Table 9: Prompt template for extracting song style (e.g. vocal type, instruments) from audio song.
@@
Task: You are a talented songwriter, music director, and composer. Your task is to compose a {genre} genre song about {topic} with a {mood} mood. Provide the lyrics and style of the song after the “Answer” section. Follow the step-by-step instructions provided in the "Instructions" section and respond using the format given in the "How to Answer" section. Instructions: 1. Before you write the song, you need to plan what you will write about the song, then synthesize the features of the song lyrics mentioned below. Subject Matter: Write what the song is about by providing a summary of the story, narrative, central topic, or events discussed in the song lyrics. Theme: Identify the main theme or message (emotion, life experience, social commentary, or philosophical concept). Target Audience: Define the intended audience (age group, cultural background, or specific interests). Narrative: The story and point of view (e.g., first person or second person). Character Analysis: Main characters (e.g., Protagonist, Antagonist) and their traits. Song Structure: Outline the structure (number of verses, choruses, bridges, intros, and outros) and note any unique elements or deviations. Mood: Describe the overall mood (upbeat, melancholic, introspective, etc.). Reference: Identify any cultural, social, time, place, or contextual references. 2. Then, you need to write song lyrics that include all the features described in the “Lyrics Features” section. The song should not be long; rather, it should be medium-length. You should also write the lyrics with metatags following the song structure from "Lyrics Features." In some very rare cases, you can also scarcely include ad libs, or non-lexical vocables. Here are the detailed instructions: You can add metatags to your lyrics on top of a section in [square brackets] that will create certain styles. Some examples of metatags are [Verse], [Chorus], [Bridge], [Solo], [Outro], [Pre-Chorus], [Bridge], [Hook], [Opening], [Intro], [Instrument], [Build], [Drop], [Breakdown], [Refrain], [Spoken], [Interlude], [Prelude], [Sample], etc. Adding a blank newline between sections yields the best results. In some very rare cases, you can also scarcely use Ad libs (vocal embellishments) to your prompts in (parentheses), only when necessary. Examples include (yeah), (alright), (come on), (whoa), etc. Ad libs tend to work best at the end of a line but can also work mid-line. Unlike metatags, ad libs are sung/verbalized. In some very rare cases, you can also scarcely use non-lexical vocables, only when necessary. Examples include la la la, na na na, sha na na. 3. Finally, you need to compose the song by synthesizing all relevant and detailed information about the song’s style. This includes identifying characteristics such as instruments used (e.g., guitar, drums, piano, violin, synthesizers, bass, orchestral, solo, acoustic, trumpet, saxophone, flute, cello, harmonica, banjo, accordion, etc.), vocals types (female or male), genres (e.g., rock, pop, pop rock, indie rock, hard rock, metal, heavy metal, r&b, electronic, soul, jazz, country, reggae, classical, hip hop, blues, folk, punk, funk, disco, alternative, grunge, etc.), tempo (slow, moderate, fast), mood (e.g., melancholic, upbeat, aggressive, melodic, sad, happy, excited, nostalgic, mellow, serene, joyful, dark, gothic, etc.), and any other stylistic elements (e.g., dance, party, cinematic, dreamy, energetic, relaxing, anthemic, atmospheric, groovy, etc.) that contribute to the overall vibe, environment, or atmosphere of the song. The name of the style must be written in lowercase and separated by commas. 4. Finally, choose a title for the song that best suits its lyrics and style. Prompt continued on next page…
@@
Table 10: Prompt template for generating song lyrics and style from genre, topic and mood. It also generates lyrics feature and song title as by-product.

### v4

@@ Appendix C Implementation Details
Appendix D Benchmark
D.1 Generalization Test
Although generalization is not the primary focus of our work, it remains crucial for evaluating the broader applicability of our dataset and methods. To assess generalization, we evaluated models trained on the SONICS dataset using out-of-distribution (OOD) songs from two external generators: SkyMusic and SeedMusic. Due to the lack of an API for SkyMusic and a platform for SeedMusic, we manually collected 394 and 36 songs, respectively, from their publicly available demo websites. While the sample size is limited, these tests offer preliminary insights into generalization performance. The results, summarized in Table 8, reveal the following key observations:
  1. 1.
    Performance Decline: A consistent decline in performance is evident across all models, as shown in Table 8. This trend aligns with findings in media forensics Ojha et al. (2023); Epstein et al. (2023); Yan et al. (2024), underscoring the challenges of generalization. For instance, ConvNeXt, which performed exceptionally well on the SONICS dataset, exhibited significant performance drops, ranking the lowest across both sources and durations.
  2. 2.
    SpecTTTra’s Robustness: SpecTTTra-$\gamma$ achieved the highest F1 scores (80% on SeedMusic and 60% on SkyMusic), outperforming all other models. However, larger variants, such as SpecTTTra-$\alpha$, showed susceptibility to overfitting, particularly with long-duration songs.
  3. 3.
    Transformer Advantage: Models incorporating transformers (e.g., SpecTTTra, ViT, EfficientViT) demonstrated greater resilience to OOD songs compared to CNN-based models, reaffirming their suitability for addressing generalization challenges.
The observed performance declines on OOD songs are consistent with prior findings Ojha et al. (2023); Epstein et al. (2023); Yan et al. (2024), emphasizing the necessity of advanced generalization strategies. Despite the absence of such strategies, some models achieved promising OOD results, highlighting the potential of the SONICS dataset for advancing song forensics. Notably, for validation purposes, we also evaluated all models with randomly initialized weights. None of these models exceeded an F1 score of 0.51, further reinforcing the significance of learned representations.
Table 8: Comparison of SpecTTTra and Existing Methods on Unseen Generators (SkyMusic and SeedMusic). AASIST and its variants face out-of-memory error while training for 120s audio samples resulting in missing performance scores (-).
Len.
Model Seed Music Sky Music
5 AASIST 0.65 0.56
ResNet + Spec. 0.52 0.48
ResNet + LFCC 0.54 0.51
Wav2Vec2 + AASIST 0.69 0.53
ConvNeXt 0.28 0.21
ViT 0.68 0.57
EfficientViT 0.67 0.54
SpecTTTra-$\gamma$ 0.32 0.33
SpecTTTra-$\beta$ 0.32 0.41
SpecTTTra-$\alpha$ 0.47 0.57
120 AASIST - -
ResNet + Spec. 0.53 0.35
ResNet + LFCC 0.56 0.40
Wav2Vec2 + AASIST - -
ConvNeXt 0.36 0.19
ViT
0.64 0.53
EfficientViT 0.71 0.55
SpecTTTra-$\gamma$ $\underline{\mathbf{0.80}}$ $\underline{\mathbf{0.60}}$
SpecTTTra-$\beta$ 0.53 0.28
SpecTTTra-$\alpha$ 0.56 0.25
D.2 Human-AI Benchmark
To assess human performance in synthetic song detection, we developed a Huggingface space called “Song Arena” as illustrated in Fig. 11. In this space, users can evaluate whether a randomly selected song from a subset of the proposed dataset (comprising 520 samples) is synthetic or not. The space also features a leaderboard (shown in Fig. 11) that records human detection performance for songs generated by different algorithms and generation methods. The evaluation metrics used to assess the detectability of synthetic songs include the F1 score, Sensitivity (True Positive Rate), and Specificity (True Negative Rate). These metrics provide a comprehensive measure of the difficulty humans face in detecting synthetic songs generated by various algorithms.
Figure 10: Song Arena on Huggingface.
Figure 11: Song Arena Leaderboard on Huggingface.
Appendix E Result Analysis
Our analysis of the SpecTTTra-$\alpha$ model’s results reveals audible and perceptible artifacts in both successful and failed cases of real and fake songs. In True Negative cases, we find distinct patterns in correctly classified real songs. These include characteristics such as unpredictability, dynamic variation, and unexpected changes that is often absent in fake songs. Examples include non-standard pitch variations, intricate rhythmic complexity, and expressive techniques like melismatic phrasing, sudden tempo changes, or improvisational segments, all of which showcase the nuanced artistry of human performance. Conversely, in True Positive cases, we detect specific audible artifacts in correctly classified fake songs. These artifacts include mechanical or robotic vocal qualities, unclear vocal articulation, predictable rhythmic structures, and limited pitch variability. Such fake songs also lack the emotive expressiveness and complexity we consistently find in real music, making them notably distinct. In False Negative cases, we observe that fake songs not detected as such lack the typical artifacts seen in true positive fake songs. Instead, these cases often incorporate features that mimic the unpredictability and nuanced variation of real songs. For instance, some include spoken interludes or conversational segments before the singing starts, creating a deceptive resemblance to genuine music. In False Positive cases, we find that real songs misclassified as fake share characteristics with AI-generated music. These include unclear vocals, less rhythmic complexity, and a noticeable absence of the unexpected changes that typically distinguishes human-created performances. Finally, we encounter instances where we are unable to detect any audible artifacts. This suggests the presence of subtle, imperceptible, or inaudible artifacts (Barrington et al., 2023) akin to invisible artifacts (Chhabra et al., 2023) in synthetic images.
Appendix F Comparison with Related Works
To comprehensively evaluate our proposed method, we compared our SpecTTTra models with various related works in fake song detection and provided our findings in Table 9. Specifically, we benchmarked our method against all the approaches mentioned in SingFake Zang et al. (2024b), including AASIST Jung et al. (2022), Wav2Vec Baevski et al. (2020), and ResNet He et al. (2016) variants. Our analysis demonstrates that the SpecTTTra models outperform the compared methods in terms of effectiveness while maintaining efficiency. For example, although the AASIST models deliver promising results for short songs, they are computationally intensive, as reflected by their high time and memory consumption. Additionally, AASIST models fail to process long songs of 120 seconds, resulting in out-of-memory (OOM) errors due to its extreme computational complexity. On the other hand, while ResNet variants exhibit high efficiency, they lack the capacity to detect fake songs as effectively as SpecTTTra, particularly with long-duration songs. In contrast, the SpecTTTra models strike a balance between efficiency and effectiveness. As shown in the Table 9, SpecTTTra-$\alpha$ achieves the highest F1 score of 0.97 for long-duration songs while maintaining manageable computational requirements. These results establish the superiority of our method in both detecting fake songs and handling longer audio inputs compared to other existing methods.
Table 9: Performance and efficiency comparison with existing works in fake song detection. AASIST and its variants face out-of-memory error while training for 120s audio samples resulting in missing performance scores (-). Their FLOPs and activation count were also unable to be determined by the fvcore library.
Len.
Model
Speed
(A/S) $\uparrow$
FLOPs
(G) $\downarrow$
Mem.
(GB) $\downarrow$
# Act.
(M) $\downarrow$
# Param.
(M) $\downarrow$
F1 $\uparrow$
5 AASIST 55 12 6 96 0.3 0.91
ResNet + Spec. 354 0.6 0.1 0.8 11 0.86
ResNet + LFCC 331 0.6 0.1 0.8 11 0.88
Wav2Vec2 + AASIST 42 * 1.3 * 95 0.90
SpecTTTra-$\gamma$ 154 0.7 0.1 2 17 0.76
SpecTTTra-$\beta$ 152 1.1 0.2 2 17 0.78
SpecTTTra-$\alpha$ 148 2.9 0.5 6 17 0.80
120 AASIST 2 295 OOM 2393 0.3 -
ResNet + Spec. 146 17.2 2.3 24 11 0.90
ResNet + LFCC 144 17.2 2.3 25 11 0.91
Wav2Vec2 + AASIST 2 * OOM * 95 -
SpecTTTra-$\gamma$ 97 10.1 1.6 20 24 0.88
SpecTTTra-$\beta$ 80 14.0 2.3 29 21 0.92
SpecTTTra-$\alpha$ 47 23.7 3.9 50 19 0.97
33footnotetext: This is a footnote explaining something about the table.
Appendix G Limitations and Future Work
The proposed SONICS dataset contains real songs dynamically queried from YouTube using their titles and artist names, which can sometimes result in incorrect audio retrieval. A manual analysis of 600 random samples suggests that this issue affects approximately 0.5% of the dataset. To address this minor noise, we utilized label smoothing (Szegedy et al., 2016) during training. Another limitation is that the fake songs generated by the Udio platform cannot include lyrics from real songs, limiting the comprehensive evaluation in the Half Fake songs category to only those generated by the Suno platform. Our current benchmarks are based solely on Mel Spectrogram inputs; hence, we aim to incorporate raw audio and explore other feature extraction methods, such as LFCC and MFCC, to enhance the robustness of our evaluations. Due to resource constraints, we trained and compared only smaller versions of all models. In the future, we plan to compare larger versions of all models. Furthermore, we trained all models from scratch to ensure a fair comparison, because the proposed SpecTTTra model is designed specifically for audio, while other models like ConvNeXt and ViT only have pretrained weights available for images (ImageNet). In the future, we plan to pretrain all models on an large-scale audio dataset from scratch before training them on our proposed dataset for maximizing the performance.
Appendix H Prompt Engineering
Selection of LLMs: For lyrics and song-style generation, as well as lyrics feature extraction, we evaluated several proprietary LLMs, including GPT-4o, Claude 3, and Gemini 1.5, along with open-source models like LLama 3, Gemma 2, and Mistral Large. Among these models, GPT-4o demonstrated superior performance, particularly in maintaining rhythm and coherence and accurately following the content of prompts. Based on these qualities, we selected GPT-4o for both lyrics and song-style generation tasks.
For song-style analysis, only the proprietary models Gemini 1.5 Pro and Gemini 1.5 Flash, as well as the open-source model Qwen-Audio (Chu et al., 2023), are equipped to process audio inputs. Our evaluations indicated that both Gemini 1.5 Pro and Gemini 1.5 Flash models deliver similarly accurate performance. In contrast, Qwen-Audio frequently struggled to follow instructions correctly. Considering that the Gemini 1.5 Flash model is ten times more cost-effective than the Gemini 1.5 Pro, we selected the Gemini 1.5 Flash model for song-style analysis.
@@ Appendix H Prompt Engineering
Instructions:
Analyze the provided song lyrics and extract the following elements:
1. Subject Matter: Write what the song is about by providing a summary of the story, narrative, central topic, or events discussed in the song lyrics.
2. Theme: Identify the main theme or message (emotion, life experience, social commentary, or philosophical concept).
3. Target Audience: Define the intended audience (age group, cultural background, or specific interests).
4. Narrative: The story and point of view (e.g., first person or second person).
5. Character Analysis: Main characters (e.g., Protagonist, Antagonist) and their traits.
6. Song Structure: Outline the structure (number of verses, choruses, bridges, intros, outros) and note any unique elements or deviations.
7. Mood: Describe the overall mood (upbeat, melancholic, introspective, etc.).
8. Reference: Identify any cultural, social, time, place, or contextual references.
How to respond:
You should provide your answer below after the “Answer” section. You are not allowed to use any text formatting (bold, italic, etc.) and narrative (‘Here is the answer’, ‘Below is the response’, etc.) in your answer. Only answer using the following format:
subject_matter: “…..”
theme: “….”
target_audience: “….”
narrative: “….”
character_analysis: “….”
song_structure: “….”
mood: “….”
reference: “….”
Lyrics:
{lyrics}
Answer:
Table 10: Prompt template for extracting lyrics features (e.g subject matter, theme, mood) from real songs. Here, {lyrics} indicates placeholder for input lyrics.
@@
You should write the song with metatags following the song structure from “Lyrics Features”. In some very rare cases, you can also scarcely include ad libs, or non-lexical vocables.
@@
Table 11: Prompt template for generating song lyrics from lyrics features. Here, {lyrics_feature} indicates placeholder for input lyrics features.
@@
Table 12: Prompt template for extracting song style (e.g. vocal type, instruments) from audio song. Here, {song} indicates placeholder for input audio song.
@@
Task: You are a talented songwriter, music director, and composer. Your task is to compose a {genre} genre song about {topic} with a {mood} mood. Provide the lyrics and style of the song after the “Answer” section. Follow the step-by-step instructions provided in the “Instructions” section and respond using the format given in the “How to Answer” section. Instructions: 1. Before you write the song, you need to plan what you will write about the song, then synthesize the features of the song lyrics mentioned below. Subject Matter: Write what the song is about by providing a summary of the story, narrative, central topic, or events discussed in the song lyrics. Theme: Identify the main theme or message (emotion, life experience, social commentary, or philosophical concept). Target Audience: Define the intended audience (age group, cultural background, or specific interests). Narrative: The story and point of view (e.g., first person or second person). Character Analysis: Main characters (e.g., Protagonist, Antagonist) and their traits. Song Structure: Outline the structure (number of verses, choruses, bridges, intros, and outros) and note any unique elements or deviations. Mood: Describe the overall mood (upbeat, melancholic, introspective, etc.). Reference: Identify any cultural, social, time, place, or contextual references. 2. Then, you need to write song lyrics that include all the features described in the “Lyrics Features” section. The song should not be long; rather, it should be medium-length. You should also write the lyrics with metatags following the song structure from “Lyrics Features.” In some very rare cases, you can also scarcely include ad libs, or non-lexical vocables. Here are the detailed instructions: You can add metatags to your lyrics on top of a section in [square brackets] that will create certain styles. Some examples of metatags are [Verse], [Chorus], [Bridge], [Solo], [Outro], [Pre-Chorus], [Bridge], [Hook], [Opening], [Intro], [Instrument], [Build], [Drop], [Breakdown], [Refrain], [Spoken], [Interlude], [Prelude], [Sample], etc. Adding a blank newline between sections yields the best results. In some very rare cases, you can also scarcely use Ad libs (vocal embellishments) to your prompts in (parentheses), only when necessary. Examples include (yeah), (alright), (come on), (whoa), etc. Ad libs tend to work best at the end of a line but can also work mid-line. Unlike metatags, ad libs are sung/verbalized. In some very rare cases, you can also scarcely use non-lexical vocables, only when necessary. Examples include la la la, na na na, sha na na. 3. Finally, you need to compose the song by synthesizing all relevant and detailed information about the song’s style. This includes identifying characteristics such as instruments used (e.g., guitar, drums, piano, violin, synthesizers, bass, orchestral, solo, acoustic, trumpet, saxophone, flute, cello, harmonica, banjo, accordion, etc.), vocals types (female or male), genres (e.g., rock, pop, pop rock, indie rock, hard rock, metal, heavy metal, r&b, electronic, soul, jazz, country, reggae, classical, hip hop, blues, folk, punk, funk, disco, alternative, grunge, etc.), tempo (slow, moderate, fast), mood (e.g., melancholic, upbeat, aggressive, melodic, sad, happy, excited, nostalgic, mellow, serene, joyful, dark, gothic, etc.), and any other stylistic elements (e.g., dance, party, cinematic, dreamy, energetic, relaxing, anthemic, atmospheric, groovy, etc.) that contribute to the overall vibe, environment, or atmosphere of the song. The name of the style must be written in lowercase and separated by commas. 4. Finally, choose a title for the song that best suits its lyrics and style. Prompt continued on next page…
@@
Table 13: Prompt template for generating song lyrics and style from genre, topic and mood. It also generates lyrics feature and song title as by-product. Here, {genre}, {topic}, {mood} indicates placeholders for input genre, topic and mood of the song.